OpenStudy (anonymous):

Essay; show all work. Last fall, a gardener planted 64 iris bulbs. She found that only 56 of the bulbs bloomed in the spring. Find the empirical probability that an iris bulb of this type will bloom. Give answer as a fraction in lowest terms. How many of the bulbs should she plant next fall if she would like at least 59 to bloom?

OpenStudy (anonymous):

\[\frac{56}{64}=\frac{7}{8}\]

OpenStudy (anonymous):

second question is ambiguous but your teacher probably wants you to solve \[\frac{59}{x}=\frac{7}{8}\] for x

OpenStudy (anonymous):

answer would be \[x=\frac{59\times 8}{7}=67.42\] so i would say 68

OpenStudy (anonymous):

how did u get 59/x=7/8?

OpenStudy (anonymous):

oh you want 59 to bloom yes? 64 gives you 56 blooms so the ratio of "blooms" to "planted" is 7 to 8

OpenStudy (anonymous):

you want 59 to something = 7 to 8, so you solve this via \[\frac{59}{something} = \frac{7}{8}\]

OpenStudy (anonymous):

oh ok...

OpenStudy (anonymous):

if not clear we can think of it this way: \[\frac{7}{8}=.875\] so 87.5% of the plants bloom. then it is like solving a percent problem "59 is 87.5% of what number?" the solution is the same

OpenStudy (anonymous):

put in lots of words so teacher is impressed with your essay!

OpenStudy (anonymous):

oh ok.. that makes sense

OpenStudy (anonymous):

heres another one...

OpenStudy (anonymous):

Essay; show all work. A community 5K run will award $50 to the winner. 65 people enter the race, and they each pay an entry fee of $20. Assuming they are all equally likely to win, what is a fair price for the competition? Round to the nearest cent.

OpenStudy (anonymous):

wouldnt you take 50-20

OpenStudy (anonymous):

i am a bit confused. it says the entry fee is $20. is that what you are supposed to adjust?

OpenStudy (anonymous):

or are you supposed to adjust the winning amount?

OpenStudy (anonymous):

i have no clue this is the question they asked.. and thats it.

OpenStudy (anonymous):

well if they are only awarding $50 to the winner then each entrant should only pay \[\frac{50}{65}\] dollars and that would break even

OpenStudy (anonymous):

if they each payed 77 cents (rounded) then that would be $50 total. since each is equally likely to win they should each pay the same amount.

OpenStudy (anonymous):

it asked for fair price for the competition round to nearest cent

OpenStudy (anonymous):

so it would be 80 cents?

OpenStudy (anonymous):

no to the nearest cent is 77

OpenStudy (anonymous):

oh ok...

OpenStudy (anonymous):

50/65=.7692... rounded give .77 or 77 cents

OpenStudy (anonymous):

one last problem... lol ty for helping me with this...

OpenStudy (anonymous):

According to the U. S. Census Bureau, the total 2008 U.S. population was 303,824,640. The chart below summarizes the 2008 population for five U.S. States.

OpenStudy (anonymous):

U. S. State 2008 Population Missouri 5,911,605 Pennsylvania 12,448,279 Tennessee 6,214,888 Utah 2,736,424 Washington 6,549,224

OpenStudy (anonymous):

What is the probability that a randomly selected U.S. resident did not live in Missouri? Show step by step work. Round solution to the nearest thousandth.